# Creative and geometric times in physics, mathematics, logic, and philosophy

### v1

@@ III Two times in Mathematics
Although it is not common to discuss time in mathematics at all, the problem of the two different kinds of time introduced above has almost a perfect counterpart in mathematics. Let us now consider a standard mathematical object, e.g. a real number. Typical real numbers – in fact, almost all of them in a mathematical sense – are uncomputable, i.e. for each of these numbers there exists no algorithm – which is by definition a finite list of instructions – that outputs all their digits (they have infinite Kolmogorov complexity) [10]. On the contrary, all the (irrational) numbers that we usually consider as prototypical examples of reals, such as $\sqrt{2}$ or $\pi$, are all computable (the ratio between the diagonal of a square to its edge is a simple algorithm that outputs all the digits of $\sqrt{2}$ and similarly the ratio of the circumference to the diameter of a circle for $\pi$). Obviously, all the rational numbers are also computable: they can be directly written down or compressed (in the case of repeating decimals) in a finite string.
@@
The standard way of regarding mathematics, the so-called classical-mathematics, is a form of Platonism which posits that, among the other mathematical entities, real numbers exist with their infinite series of digits, although there is in general no way to even label and thus grasp them (since there is no algorithm that generates them [20, 21]. Hence, classical Platonistic mathematics is a timeless language, which per se is not a bug, but it becomes problematic when mathematics is elevated to the language of science, i.e., it is used to the describe the physical world.
@@
However, there are alternative approaches known as constructive mathematics, of which the most prominent is intuitionism. Therein, mathematical entities are not given at once, but rather are processes in development (digits get created continuously, one after the other). This provides mathematics with a concept of passage of time (and it would be only at infinite time – at the end of time, so to say – that the mathematical entities, such as real numbers, are completed into the ones defined by classical mathematics). The initiator of intuitionism, L.E.J. Brouwer, envisioned a ``````` `creative subject”, i.e., an idealized mind or mathematician, who is responsible for this progressive process of creation in time of, for instance, the digits of an uncomputable real numbers. Several authors, however, have distanced themselves from this controversial concept of a creative subject. In particular, one of us (N.G.) has put forward the idea that the digits of a typical real number are generated by a true random number generator, [^4] More precisely, it is not necessarily new digits that come into existence, but the new information reduces the indeterminacy of intuitionistic real numbers [21]. i.e., a natural process that is able to create a piece of new information by changing a fundamentally indeterminate bit into a determined one [10]. [^5] For more works that relate constructive mathematics to physics, see [20] and [22]. This version of intuitionism could thus be labeled ``````` `objective” or ``````` `naturalistic” intuitionism. In this way, typical real numbers become a graspable concept directly linked to creative time, that is, to the change from the indeterminate to the determinate. Hence, naturalistic intuitionism, is a tensed mathematical language.
@@
To think about computable real numbers, one just needs geometric time. In fact, it is possible to think that the full information about those numbers is contained in the (finite) algorithm that defines them, i.e., in their initial conditions. For instance, one can ask what is the 43800th digit of $\pi$ and the answer is given by running an algorithm that outputs the digits of $\pi$ and picking its 43800th digit. Note that it is not necessary to go through all the previous digits of $\pi$ to compute 43800th one. There are so-called digit-extraction algorithms that allow one to directly compute the nth digit of $\pi$ [23]. [^6] The most well-known of such algorithms is perhaps the Bailey-Borwein-Plouffe formula [24], that allows to directly compute the hexadecimal digits of $\pi$. Digit-extraction algorithms, also in base 10, are known for several other irrational computable real numbers, such as for $e$. This can be seen as a further evidence that computable (irrational) numbers are already fully determined and do not require creative time. The point is that since this is found through a deterministic outcome, the answer to this kind of questions is fully contained in the algorithm, which can be run more or less fast, giving the ability to manipulate this ``````` `mathematical geometric time”. This exactly resembles the weather example in Sect. II for the deterministic evolution of the physical system that forms the weather.
@@
In contrast, if one considers instead a typical, i.e., uncomputable, real number, things are different. There are no algorithms compressing the information of that number. The only existing ``````` `algorithm” is the number itself. Each next digit is generated by a genuinely natural random process, therefore, asking which is the 43800th digit of such a number requires to wait for all 43800 instances of creating of the ``````` `next digit” (see footnote 4), i.e., it requires ``````` `mathematical creative time”.
@@
Finally, one can also ask a question of the kind ``````` `are there 700 consecutive sevens in the digits of a number. [^7] The example of 700 consecutive sevens is borrowed from C. Posy [25]. If that number is uncomputable, then one has no choice but to wait and see. However, even for computable numbers, like e.g. $\pi$, the question is interesting. The situation is similar to the example of the weather – i.e., 700 consecutive days of rain if one assumes determinism – but the axiomatic character of mathematics makes the solution to this question clearer. One can, in fact, think of programming two softwares: The first, $S_{1}$ is exactly the one we used in the physical example of the weather, i.e., the systematic search of a 700 consecutive sevens in $\pi$; on the other hand, the second software, $S_{2}$ outputs all the theorems that can be derived within Peano arithmetic and it halts should it find the negation of the conjecture. If the statement is true, $S_{1}$ will find this sequence of 700 sevens and therefore halt after a finite (geometric) time. If the statement is provable (within Peano’s arithmetic), $S_{2}$ will halt after a finite (though presumably enormous) geometric time. Interestingly, if one assumes that neither halts, then the statement is false (if not $S1$ would halt), but not provable within Peano’s arithmetic (if not $S_{2}$ would halt). [^8] For the sake of completeness, there is also the fourth logical possibility that both $S_{1}$ and $S_{2}$ halt, a case which would prove Peano’s arithmetic inconsistent. This would be an example of Gödel’s celebrated theorem. Note that mathematicians don’t use softwares like $S_{1}$ and $S_{2}$, but use their creativity to find shortcuts to analyse such statements, i.e., de facto they use creative time. In the case of the weather in Geneva, if one assumes that the system is indeterministic, like uncomputable numbers, then there is no choice but wait and see. However, if one trusts the equations describing the weather evolution and the fully determined initial conditions, then one could run analogues of $S_{1}$ and $S_{2}$, as for computable numbers. Note however, that this assumes one can compress the initial conditions into a finite algorithm (hence into a computer which is necessarily finite).
@@
Let us acknowledge that this distinction between geometric and creative time in mathematics distances our view from the most accepted position in intuitionism, such as that of C. Posy [25] or M. van Atten [26, 27], who consider all numbers, whether computable or not, as generated by new information that comes about, so that they all require creative time in our parlance.
@@ IV Logic
[…]
Figure 2: Representation of ontic indeterminacy in phase space.
Now, the principle of non-contradiction imposes already that propositions about physics must be referred to a time. Indeed, if one states simpliciter that ``````` `the system is inside $r$” and ``````` `the system is not inside $r$”, this is a violation of the principle of non-contradiction, since we do observe in nature the same system being sometimes inside $r$ and sometimes not. One has then to refer to a time to qualify the propositions. If one adds ``````` `at time $t$” to both propositions, then the principle of non-contradiction holds because –at any time $t$ – only one of these propositions would be true and the other not. On the contrary, if one labels one proposition with ``````` `at time $t_{1}$” and the other with ``````` `at time $t_{2}$”, with $t_{1}\neq t_{2}$, it is possible that they are both true. Hence, the principle of non-contradiction applies only to propositions at the same time. Note that this time is not necessarily creative, but it could just be the label appearing in the equations of motion, i.e., geometric time.
Coming instead to a scenario where there is ontic indeterminacy, we now show that the law of the excluded middle fails. Consider for instance a model of physics (classical, quantum, or even post-quantum) in which a (pure) state is represented by a finite volume, $R$, in phase-space (as opposed to a dimensionless mathematical point). That is, there is not merely a lack of knowledge about the system (as it is the case in, e.g., statistical physics), but this state contains the complete information existing at present about all the properties of the system (see also [3, 4]). Since, as we have already argued, this ontic indeterminacy generally leads to indeterminism in the future evolution of a system, [^10] Except for integrable classical dynamical systems. statements about future properties are in general neither true nor false because they are yet to be determined. And it is this process of determination that requires creative time. In general, in the case of ontic indeterminacy, one needs to introduce a multivalued tensed logic according to which a proposition is either true or false only at a time $\tau\geq\tau_{0}$ when it has become determined, but its truth value was ``````` `indeterminate” (as a third logical value) before $\tau_{0}$ (for multivalued logic see [28, 29]). As an example, consider a true random number generator that outputs a (single determined) bit of information, say $b=0$. That event defines the (creative) time instant $\tau_{0}$. After such an event that brings into existence this new bit of information, the truth value of the proposition ``````` `b=0” is true and the proposition ``````` `b=1” is false, but before $\tau_{0}$ both propositions had a truth value ``````` `indeterminate”. This represents a breakdown of the principle of the excluded middle. Note that such a failure of the excluded middle is found in all constructive mathematics – e.g. intuitionism [25]. However, despite these many similarities, we emphasize that our approach differs substantially from ``````` `orthodox” intuitionism. First of all, as already remarked in Section III, we reject the Brouwerian notion of ``````` `creative subject” in intuitionistic logic, advocating for a naturalistic version of intuitionism in which new information is created in the universe by natural processes. Moreover, contrarily to what we have advocated, we stress that orthodox intuitionistic logic is not three-valued (nor multivalued for that matter) – as proven by Gödel [30]– and that logical propositions are typically timeless.
Finally, consider a system whose present state is represented by a finite volume in phase space $R$ (see Fig. 2). This means that although the state is fully determined (but not as a collection of real numbers [9]), it does not fully determine the future. In this case, a statement like ``````` `the system is located in phase space within $r\subset R$” may lead to think that the principle of the excluded middle fails also for statements about the present. However, such a statement is not indeterminate, but actually false, for it is simply not the case that the system is located inside $r$ in phase space. The only presently true statement about the state of the system is ``````` `the system is located at $R$ in phase space” and nothing more. On the contrary, the statement ``````` `the system will be (at some future time) located in phase space at $r$” is currently indeterminate, because it is not impossible that the system will become more localized (in phase space). Once the evolution of the system develops, the statement will become either true or false. This process is what requires creative time. Hence, the failure of the excluded middle concerns only future statements, where the future is identified by a change of creative time.

### v2

@@ III Two times in Mathematics
Although it is not common to discuss time in mathematics at all, the problem of the two different kinds of time introduced above has almost a perfect counterpart in mathematics. Let us now consider a standard mathematical object, e.g. a real number. Typical real numbers – in fact, almost all of them in a mathematical sense – are uncomputable, i.e. for each of these numbers there exists no algorithm – which is by definition a finite list of instructions – that outputs all their digits (they have infinite Kolmogorov complexity) [48]. On the contrary, the (irrational) numbers that we usually consider as prototypical examples of reals, such as $\sqrt{2}$ or $\pi$, are all computable (the ratio between the diagonal of a square to its edge is a simple algorithm that outputs all the digits of $\sqrt{2}$ and similarly the ratio of the circumference to the diameter of a circle for $\pi$). Obviously, all the rational numbers are also computable: they can be directly written down or compressed (e.g., in the case of repeating decimals) in a finite string.
@@
The standard way of regarding mathematics, the so-called classical-mathematics, is a form of Platonism which posits that, among the other mathematical entities, real numbers exist with their infinite series of digits, although there is in general no way to even label and thus grasp them (since there is no algorithm that generates them [45, 12]). Hence, classical Platonistic mathematics is a timeless language, which per se is not a bug, but it becomes problematic when mathematics is elevated to the language of science, i.e., is used to describe the physical world.
@@
However, there are alternative approaches known as constructive mathematics, of which one of the most prominent is intuitionism (see [15, 73, 10] and references therein). Therein, mathematical entities are not given at once, but rather are processes that develop in time (digits get created continuously, one after the other). [^6] More precisely, it is not necessarily new digits that come into existence, but the new information reduces the indeterminacy of intuitionistic real numbers [12]. This explicitly provides mathematics with a concept of passage of time (and it would be only at infinite time – at the end of time, so to say [73] – that the mathematical entities, such as real numbers, are completed into the ones defined by classical mathematics). Indeed, one should stress that “the dependence of intuitionism on time is essential: statements can become provable in the course of time and therefore might become intuitionistically valid while not having been so before”. [50].
The initiator of intuitionism, L.E.J. Brouwer, envisioned intuitionistic mathematics as “having its origin in the perception of a move of time” [16]. However, for him mathematics was the activity of the mind of a ``creative subject”, i.e., an idealized mind or mathematician, who is responsible for this progressive process of creation in time of, for instance, the digits of an uncomputable real numbers [16]. Note that, for Brouwer, time is merely internal, i.e., it is not physical time at all – neither geometric nor creative in the sense discussed here (see [9]). Several authors, however, have distanced themselves from this controversial concept of a creative subject. In particular, one of us (N.G.) has put forward the idea that the digits of a typical real number are generated by a “natural” random number generator (see also footnote 6), i.e., a natural process – as opposed to the action of an agent – that creates new information in the world by changing a fundamentally indeterminate bit into a determinate one [47]. [^7] For more works that relate constructive mathematics to physics, see [45] and [42]. This version of intuitionism, which departs from the standard view (see e.g., [73, 86]), could thus be labeled ``objective” or ``naturalistic” intuitionism, as characterized in [8]. In this way, typical real numbers become a graspable concept directly linked to creative time, that is, to the change from the indeterminate to the determinate happening in natural phenomena. Hence, naturalistic intuitionism, is a tensed mathematical language.
@@
To think about computable real numbers, one just needs geometric time. In fact, it is possible to think that the full information about those numbers is contained in the (finite) algorithm that defines them, i.e., in their initial conditions. For instance, one can ask what is the 43800th decimal place of $\pi$ and the answer is given by running an algorithm that outputs the digits of $\pi$ and picking its 43800th decimal. The answer happens to be a 9, and it is easily computable because it is not necessary to go through all the previous digits of $\pi$ to compute 43800th one. There are so-called digit-extraction algorithms that allow one to directly compute the nth digit of $\pi$ [70]. [^8] The most well-known of such algorithms is perhaps the Bailey-Borwein-Plouffe formula [5], that allows to directly compute the hexadecimal digits of $\pi$. Digit-extraction algorithms, also in base 10, are known for several other irrational computable real numbers, such as for $e$. This can be seen as a further evidence that computable (though irrational) numbers are already fully determinate and do not require creative time. The point is that since this is found through a deterministic outcome, the answer to this kind of question is fully contained in the algorithm, which can be run more or less fast, giving the ability to manipulate this ``mathematical geometric time”. This exactly resembles the weather example in Sect. II for the deterministic evolution of the physical system that forms the weather.
@@
In contrast, if one considers instead a typical, i.e., uncomputable, real number, things are different. There are no algorithms compressing the information of that number. The only existing ``algorithm” is the number itself (more precisely there is no finite list of operations that generates the number, but the only way to indicate the number is to list its infinite digits). Each next digit is generated by a genuinely random natural process. Therefore, asking what the 43800th digit of such a number requires waiting through all 43800 instances of the creation of the ``next digit” (see footnote 6); that is, it requires ``mathematical creative time”.
In intuitionism, these numbers are generated in time by what are usually called non-lawlike choice sequences, i.e., sequences of computable numbers with new bits (or digits) created as time processes. [86, 73, 10, 64, 65]. [^9] In what follows, we will consider that uncomputable numbers are generated by non-lawlike sequences, in contrast to computable numbers that are generated by lawlike sequences. In our naturalistic version of intuitionism [8], fresh bits get generated by a natural random number generator, i.e., a process occurring in nature that is able to produce true randomness, and a choice sequence $\alpha(n)$ is the outcome of a computable function that takes as an argument $\alpha(n-1)$ – where $n$ is the time index – and all the random bits previously generated (assumed to be independent and identically distributed). [^10] Note that although the computable function deterministically produces outcomes given suitable inputs, the inputs themselves are random—thus rendering the output indeterminate. Typical, uncomputable real numbers can be regarded as generated through this natural process of creation of new information as creative time processes.
This has a connection to the mathematical model of indeterminate physics that was introduced in Ref. [24]. There, the standard real number are substituted by so-called finite information quantities (FIQs), which only encode the propensity of the digits of a physical quantity to actualize. FIQs can be regarded as a particular species of choice sequences, where the next generated bit is biased by taking the majority vote of the values taken by the previous $k$ bits, with $k$ being an odd integer. In this framework, the propensities emerging in FIQs stem from the inherent randomness of $50\%-50\%$ bit choices within these sequences. If a strong majority of recent $k$ bits is 1, it is likely that this trend will persist in the near future. This likelihood is captured as a propensity, constrained to rational values. (see [47] for a detailed discussion). [^11] To fully recover the comparison between physics and mathematical intuitionism, one might expect a discussion extending beyond the notion of choice sequences. The present work, however, focuses on a more foundational stage: establishing the meaningfulness of formulating physical theory using choice sequences, particularly in light of a constructive reconsideration of time.
@@
Finally, one can also ask a question of the kind ``are there 700 consecutive sevens in the digits of a number. [^12] This example is borrowed from C. Posy [73]. If that number is uncomputable, then one has no choice but to wait and see. However, even for computable numbers, like e.g. $\pi$, the question is interesting. The situation is similar to the example of the weather – i.e., whether there will be 700 consecutive days of rain if one assumes determinism. One can, in fact, think of programming two types of software: The first, $S_{1}$, is exactly the one we used in the physical example of the weather, i.e., the systematic search of 700 consecutive sevens in $\pi$; on the other hand, the second software, $S_{2}$ outputs all the theorems that can be derived within Peano arithmetic and it halts should it find the negation of the conjecture that there are 700 consecutive sevens in $\pi$. If the conjecture is true, $S_{1}$ will find such a sequence of 700 sevens and therefore halt after a finite (geometric) time. If the conjecture is provably false (within Peano’s arithmetic), $S_{2}$ will halt. Interestingly, if neither halts, then the statement is false (otherwise, $S1$ would halt), but this is unprovable within Peano’s arithmetic (otherwise, $S_{2}$ would halt). [^13] Given that the lifetime of the universe is sufficient for carrying out the calculation. Moreover, for the sake of completeness, there is also the fourth logical possibility that both $S_{1}$ and $S_{2}$ halt, a case which would prove Peano’s arithmetic inconsistent. This would be an example of Gödel’s celebrated theorem. Note that mathematicians don’t use software like $S_{1}$ and $S_{2}$, but use their creativity to find shortcuts to analyse such statements, i.e., de facto they use creative time. In the case of the weather in Geneva, if one assumes that the system is indeterministic, like uncomputable numbers, then there is no choice but wait and see. However, if one trusts the equations describing the weather evolution and the fully determinate initial conditions, then one could run analogues of $S_{1}$ and $S_{2}$, as for computable numbers. Note however, that this assumes one can compress the initial conditions into a finite algorithm (hence into a computer which is necessarily finite).
We would like to emphasize that one of the key features of intuitionism is that, while introducing time and finiteness into mathematics, it still preserves the concept of the continuum. This makes intuitionism a suitable mathematical framework for physics, as most physical theories rely on continuity (e.g., Lorentz transformations in relativity require the continuum). However, the intuitionistic continuum is “sticky” or “viscous” [73] because, at any finite time, choice sequences are still ongoing, and one cannot select a specific point from the continuum, as it is not yet determinate whether the sequence will converge to above or below or on the specific point. This is the origin of Brouwer’s famous continuity theorem that states that all total functions , i.e., functions defined for every element of its domain, are continuous [47].
@@
Let us acknowledge that this distinction between geometric and creative time in mathematics distances our view from the most accepted position in intuitionism, such as that of C. Posy [73] or M. van Atten [87, 88], who consider all numbers, whether computable or not, as generated by new information that comes about, so that they all require creative time in our parlance. Exactly how these approaches to intuitionism diverge, philosophically, is a question we have only recently begun to investigate with B. Bentzen [8] and that will require more systematic work.
@@ IV Logic
We have seen that creative time emerges, both in physics and in mathematics, from indeterminacy, i.e., from the possibility of having new information that comes into existence. Another way to investigate the role of time and indeterminacy is to consider propositions about physical properties and analyze their logical structure.
In this logical framework, one formulates physical properties as propositions; for example, $p(R)$ denotes the proposition “the system has property RR”. The aim of this approach is to assign truth values to physical propositions and to combine them using logical connectives. In particular, $p(R)$ is true, at the empirical level, if and only if an ideal (dichotomic) measurement of the corresponding property of the system would yield the outcome $R$ with certainty. [^14] By “dichotomic” we mean that the measurement can only answer, ideally, whether the system has property $R$ or not. In a classical physical theory, this criterion can be implemented by identifying $R$ with a region of phase space and checking whether the system’s state is entirely contained in $R$.
Two central principles uphold in classical logic are the law of non-contradiction, $\neg(p\wedge\neg p)$, and the law of the excluded middle (LEM), $p\vee\neg p$. We now examine how different notions of time influence the logical structure of propositions about physical properties, and how these two fundamental principles are affected in relation to underlying ontologies, whether deterministic or indeterministic.
Determinism in logic can be defined as follows: “If $p$ is true then it has always been the case that $p$ will be true”, for all propositions $p$. Indeterminism is just its negation: $\exists p$ such that $p$ is true and it has not always been the case that $p$ will be true” [84]. In what follows, we discuss two main arguments that, in our opinion, hold in logic when applied to physics: (i) the law of non-contradiction requires a (geometric) time label for both determinism and indeterminism; and (ii) the law of the excluded middle fits best with deterministic logic (and hence with geometric time), whereas indeterminism (related to creative time) fits best with the rejection of the LEM.
(i) Law of non-contradiction
Since physical states evolve in time, propositions about physical properties, such as $p(R)$ above, must reflect this temporality. In fact, if one asserts simpliciter – without any temporal qualification – that “the system has property RR” and also that “the system does not have property RR”, one is led to an apparent contradiction. For a system whose state changes in time, each statement may be correct at different instants; hence, without a time index, both $p(R)$ and $\neg p(R)$ would be true, thereby violating the principle of non-contradiction.
A natural remedy is to supply a temporal index. Once one writes “the system has property $R$ at time tt” and “the system does not have property $R$ at time tt”, the principle of non-contradiction is respected, since at any given time only one of these propositions can be true (the pure state of a system is unique at any given time). By contrast, if the propositions concern different instants, $t_{1}\neq t_{2}$, both may be true without inconsistency. Thus the principle of non-contradiction applies only to propositions referring to the same time, a point already emphasized by Kant [51].
Hence, time is necessary for logical consistency when dealing with propositions referring to physical systems. In deterministic physics, this time can simply be the parameter appearing in the equations of motion, i.e., geometric time. In indeterministic physics, by contrast, one must consider the logical structure of propositions in which creative time plays a role.
(ii) Law of the excluded middle
The law of the excluded middle (LEM) sits naturally within a logically deterministic framework, in which the truth values of all propositions are fixed and constant in time. By contrast, in a fundamentally indeterministic physics—where future contingencies are not yet determinate and become so only at a later moment of creative time—logic models in which the LEM may fail are more naturally motivated, as already noted by Aristotle (see [92], Ch. 9). In such a setting, the truth values of propositions about the future need not be determinate in the present. To capture this situation, one may adopt non-classical logical frameworks that explicitly accommodate indeterminacy.
In the remainder of this section, we discuss some of these approaches, which illustrate the failure of the LEM and provide formal frameworks aligned with indeterminacy and the associated notion of creative time. We stress that we do not intend here either to give a systematic review of these suitable logical systems, nor to advocate for any specific one. Our main aim is to give notable examples from the literature that align with our goal of introducing temporality into logic in a way that the LEM does not hold, for the description of indeterministic physics. We will now show that this can be achieved in different ways, such as by introducing temporality explicitly into the syntax (like in Prior’s tense logic where a future operator is introduced as a new logical connective), or solely into the semantics (like in Kripke models in intuitionistic logic).
A first possibility for dealing with indeterminate propositions would be to adopt a multivalued logic – with more truth values than just true and false, for example an “indeterminate” value – as in the influential work of J. Łukasiewicz [55]. This approach, however, raises other foundational difficulties, such as the controversial definition of negation, and will therefore not be discussed further here.
On the other hand, tense logic – when combined with an open-future and certain semantics – offers a way to maintain a two-valued logic while allowing a failure of the LEM for future contingents. No third truth value is introduced: propositions are always either true or false relative to a given time and a given future branch. However, in an open-future setting, propositions about the future are evaluated with respect to multiple admissible branches. As a result, a future-tense proposition may be true on some branches and false on others, and thus the LEM need not hold globally for future contingents, even though bivalence is preserved locally on each branch. [^15] Strictly speaking, tense logic does not reject the LEM at the syntactic level; the failure arises at the semantic level under branching-time (open-future) interpretations, which we assume throughout.
This perspective was first formalized by A. Prior (see [67] and references therein) within a modal framework introducing temporal operators such as the future operator $F$, where $Fp$ is read as “the proposition $p$ will be true at some future time relative to the present.” Note that although $F$ introduces temporality explicitely in the syntax, it is not necessarily the case that this leads to a violation of LEM unless one assumes an open-future semantics of a certain kind [^16] In fact, Prior initially explored systems that allow violations of the law of non-contradiction while retaining the LEM [71], but he later developed the Peircean and Ockhamist systems [76, 67], in which non-contradiction is preserved and the LEM fails specifically for future contingents under branching-time semantics.
On this note, an important development of Prior’s ideas is the so-called branching-time semantics [7, 67]. In a branching model of time, truth at the present moment is evaluated relative to a set of admissible future branches. At the present time ($t=0$), FpFp is true if and only if, for every admissible future branch $b$, there exists a future time $t_{b}>0$ along that same branch at which $p$ holds. Conversely, $\neg Fp$ is true if and only if, for every admissible future branch $b$ and for every future time $t>0$ along that branch, $p$ fails to hold. When $p$ holds along some branches but not others, neither $Fp$ nor $\neg Fp$ is true at present, and the law of excluded middle $Fp\lor\neg Fp$ fails for such future contingents, even though no third truth value is introduced.
Here we do not commit to any particular logical system; rather, we generically advocate for a two-valued logic in which the law of excluded middle fails for future contingencies, while the principle of non-contradiction is preserved. We take this approach because, in our view, it best captures the notion of creative time in an indeterministic physics (again, this is similar to Aristotle’s position [92], however motivated in naturalistic terms). Note that such a failure of the law of excluded middle also appears in constructive mathematics—e.g., intuitionism [73, 50]—making this proposal simultaneously meet our desiderata for a logic model and remain consistent with our perspective on mathematics in Sect. III. In fact, Kripke models for intuitionistic logic can be understood temporally if we view possible worlds as construction stages that track the mental activity of Brouwer’s constructing subject [91], or as time steps associated with actualization in our naturalistic view [8]. Note, however, that temporality is not made explicit in the syntax, as it is in tense logic [91].
Let us analyze in more detail the failure of the LEM in this view. In our indeterministic framework, physical properties will eventually actualize, yielding eventual bivalence: either $p$ or $\neg p$ will be true once the relevant (creative) time has occurred. However, this eventual bivalence does not imply that either $p$ or $\neg p$ is already true at the present time. In intuitionistic logic, to say that $p$ is true at $t=0$ means that $p$ can be proven at $t=0$, using the information available at that time. in the context of indeterministic physics, “can be proven” is understood as “is in fact already determinate”, as argued in greater detail in [8]. Accordingly, for the disjunction appearing in the LEM to be true, the proposition $p\vee\neg p$ must satisfy two conditions: (i) either $p$ or $\neg p$ is true, and (ii) it is determinate which of the two alternatives holds. [^17] In standard (i.e., non-naturalistic) intuitionism, this condition is formulated epistemically: $p\vee\neg p$ is true iff (i) either $p$ or $\neg p$ is true, and (ii) it is known, or has been proven, which alternative holds. Since we assume that propositions about the future are, in general, ontically indeterminate, the LEM does not hold for them at the present time. It is precisely the process of determination – which restores the LEM for present propositions – that requires creative time. It may also be useful to consider intuitionistic free logic, since it allows to deal with objects that may not yet exist, such as worlds or branches that did not yet actualized [72].
We emphasize again that our approach differs from “orthodox” intuitionism, for we reject the Brouwerian notion of the “creative subject” in intuitionistic logic, advocating instead a naturalistic version of intuitionism in which new information is created in the universe by natural processes [8]. [^18] There is historical evidence that Brouwer discussed the possibility of sequences generated by a physical process, e.g., throwing a die [89, 90]. Although this is a matter of historical debate, it seems that Brouwer rejected this possibility, “probably because it was based on the physical world” [16].
In summary, even in logic we see natural connections between different logic models and determinism versus indeterminism. We have argued that classical logic is most compatible with a deterministic worldview and aligns naturally with geometric time, whereas tense logic and (naturalistic) intuitionistic logic—where the LEM fails—provide a better fit for indeterminism and the notion of creative time.